OpenStudy (anonymous):

What is the volume of the cylinder? Use for . 188.4 cubic miles 282.6 cubic miles 345.4 cubic miles 471 cubic miles

OpenStudy (bahrom7893):

what cylinder?

OpenStudy (bahrom7893):

V = 2pi*r*h

OpenStudy (bahrom7893):

Plug in the values you have and tell me what you get.

OpenStudy (anonymous):

376.8

OpenStudy (bahrom7893):

What values did you plug in?

OpenStudy (anonymous):

i plugged in V=2*3.14*6*10

OpenStudy (bahrom7893):

Well the formula asked you for the radius, and you plugged in the diameter, 10. How would you find the radius if u plugged in the diameter? Oh and thanks to @tcarroll010 for pointing out my mistake. V = pi*r^2*h

OpenStudy (bahrom7893):

I gave you the lateral surface area formula by accident. In any case, find the radius and the height, and then plug it into the formula: V = pi*r^2*h

OpenStudy (anonymous):

Then I got 1130.4

OpenStudy (bahrom7893):

Okay can you post your work?

OpenStudy (anonymous):

i am using a calculator

OpenStudy (bahrom7893):

yea just post what you entered into the calculator.

OpenStudy (anonymous):

V = 3.14*6^2*10

OpenStudy (bahrom7893):

Again, the formula is asking for: V = pi * r^2 * h where